### v1

# Generalizing quantum dimensions: Symmetry-based classification of local pseudo-Hermitian systems and the corresponding domain walls

Abstract
We study conformal field theories (CFTs) and their classifications from a modern perspective based on the abstract algebraic formalism of symmetries or conserved charges, known as symmetry topological field theories (SymTFTs). By studying the algebraic structure of the SymTFTs in detail, we found a natural generalization of the quantum dimensions associated with (pseudo-)Hermitian systems and (non)-unitary CFTs. These generalized data of SymTFTs provide classifications of massless and massive renormalization group flows, which will describe the quantum phase transitions of the corresponding pseudo-Hermitian systems. Moreover, our discussions straightforwardly enable one to relate a general class of coset constructions or level-rank dualities to domain wall problems between topological quantum field theories (or a series of corresponding quantum phase transitions related to the Higgs mechanism). Our work provides a systematic reduction and classification of algebraic data, symmetries, for pseudo-Hermitian systems based on ideas from established mathematical fields, linear algebra and ring theory.
pacs:
73.43.Lp, 71.10.Pm
YITP-25-174
I introduction
Non-Hermitian systems and the corresponding nonequilibrium phenomena are among the central focuses in contemporary theoretical and experimental physics. The research is growing very rapidly in these decades, but the characterization of their criticalities and topological orders (TOs) has still been in development. The literature is large, so we only list several reviews in condensed matter [1, 2], a short review on the Lindblad form [3], and an earlier review[4]. For readers in high-energy physics theory, we note that many series of non-Hermitian models and corresponding nonunitary conformal field theories (CFTs) have been studied in the communities studying integrable lattice models, with a connection to ordinary differential equations and integrable model (ODE/IM) correspondence[5]. Also, nonunitary CFTs appear in the holography of de Sitter spacetime [6, 7], traversable wormhole [8]. We note several earlier works in coset CFTs[9, 10, 11] and a work pointing out the connection to the underlying quantum groups[12]. It is remarkable that this earlier research direction has already been commented on in the pioneering work of the Hatano-Nelson model [13], a landmark non-Hermitian model.
One of the most fundamental ideas in the studies on non-Hermitian systems is the similarity transformation, which maps a pseudo-Hermitian system to the Hermitian system [14, 15, 16]. Especially, when the system has the $PT$ symmetry, this has played an important role[17]. In quantum field theory, a kind of law of thermodynamics, called Zamolodchikov’s $c$-theorem[18] has been generalized to $c_{\text{eff}}$-theorem[19] for $PT$ symmetric models where $c$ is the central charge of a CFT and $c_{\text{eff}}$ is the effective central charge. The phenomenology of this theorem is simple: the effective central charge $c_{\text{eff}}$ of a system should not increase during renormalization group (RG) flow under some suitable assumptions, such as $PT$ symmetry. Hence, one may expect a dual relation between the $c_{\text{eff}}$-theorem for non-Hermitian systems and the $c$-theorem for the corresponding Hermitian systems, but the quantum field theoretic (QFT) formulation of the similarity transformation has not been studied widely. Moreover, it is known that the similarity transformation can change the locality of the system, and this is analogous to the gauge transformations[20]. Hence, at this stage, a local pseudo-Hermitian system is more treatable than the corresponding nonlocal Hermitian system in QFT  [^1] We will study the symmetry topological field theories of nonlocal Hermitian systems constructed from local pseudo-Hermitian systems by the similarity transformations elsewhere.. In this sense, characterizations of pseudo-Hermitian systems and their quantum phase have still been under development in the QFT language. We note related earlier studies[22, 23] and recent one[24]. It is also worth noting that the corresponding level-rank duality[25, 26, 27] has been studied in [28, 29, 30] with a close connection to the 3d3d3d-3d correspondence[31, 32].
On the other hand, there has been significant progress in the understanding of the classification of CFTs and the corresponding topological quantum field theories (TQFTs) based on their symmetry algebra (or the corresponding category theories). In the previous century, the group symmetries and their classification of various models in physics have been studied widely, and this framework has been generalized to ring (and higher-algebra) structures[33, 34, 35]. Roughly speaking, this paradigm is based on studies on the ring structure corresponding to the fusion rules of anyons in a system, and the corresponding classification has been studied under various names depending on the respective research communities. The renormalization group (RG) domain wall[36, 37] and the gapped domain wall[38] are the corresponding relatively well-established concepts in high-energy physics and condensed matter, respectively. The phenomenology itself is simple. The anyons and the corresponding symmetry operators (or conserved quantities) should form a ring over the complex number field $\mathbb{C}$, and their relation can be transformed (or deformed smoothly) under the ring homomorphism. The expression of the categorical objects as a ring over $\mathbb{C}$, or linear operators in linear algebra, has appeared in the pioneering work [39]. In [39], the precise and concrete relationship between the conserved charges and topological defects satisfying the same fusion rule has been clarified. In other words, the fusion categorical structure, which had been studied in (pure) mathematical literature[40, 41, 42, 43], has become a physical structure by identifying the corresponding objects as defects or conserved charges. To emphasize the role of the set of conserved charges as a ring over $\mathbb{C}$, we denote it as fusion ring symmetry rather than fusion category symmetry in the literature[44, 45], because fusion category symmetry sometimes cannot be identified as a ring over $\mathbb{C}$ in literature.
We stress that the fusion ring, a ring over $\mathbb{C}$, is more general than the fusion category respecting nonnegative integer matrix representation (NIM-rep) at the present stage [^2] In this sense, a part of category theories appearing in physics is less abstract and less general than linear algebra.. Hence, as we will demonstrate, there can exist ring homomorphisms that can be treated in linear algebra but cannot be in such a restricted category theory. This aspect has been found in [47] and confirmed more recently in [48]. The scope of these works is conserved charges of CFTs or the corresponding excitation operators, anyons, of the corresponding topological quantum field theories (TQFTs)[49]. They are different from defects or boundary conditions respecting the integer coefficients before objects, whereas there exists the correspondence between the NIM-rep of the fusion coefficients of such operators and that of defects [^3] We thank Yuji Tachikawa and Jurgen Fuchs for the related discussions and many useful comments.. To some extent, this undesired elementary fact (or puzzle) is a reason why this simple understanding, classifications of anyons by ring homomorphisms, has not been studied widely. It should be stressed that, whereas the NIM-rep plays a significant role in studying defects, the ring over $\mathbb{C}$ is more elementary, fundamental, and historically established in studying the algebra of conserved charge in quantum Hamiltonian systems in QFTs and lattice models. For readers interested in the RG flows of defects or boundary conditions under bulk perturbations, we note earlier works on the corresponding bulk and boundary RG flows[51, 52, 53, 54] and recent works on translational invariant defects [55, 56]. They are not in the main scope of the present manuscript, but they contain interesting findings worth for further studies.
The readers who are rather familiar with recent applications of a particular (or too restricted in the above contexts) type of category theories in physics may not be familiar with treating a set of conserved charges as a ring over $\mathbb{C}$. However, this ring-theoretic or linear algebraic formulation of symmetries (or conserved charges) is more standard in other established wider research fields, such as experimental and theoretical physics, especially in quantum physics (also in a wide field of mathematics). One can see many applications of a ring over the number field, or linear algebra, ubiquitously but the class of category theories used in physics cannot capture the properties of such an established structure. Hence, the classification of the rings, such as ring isomorphism, subring, and ring homomorphism, defines the classification of theories more generally. A massive RG corresponds to an operation taking a subring, and a massless RG[57] corresponds to a ring homomorphism. The former can be understood by relating the massive RG as the addition of perturbations which break or reduce symmetry, and the latter has been proposed in the studies of RG domain walls[58, 59, 60]. For the massive RGs, to some extent, it has already been established that the ground states or the consequence of the RG can be analyzed by the corresponding (smeared) boundary conformal field theories (BCFTs)[61, 62, 63, 64, 65, 66, 67, 68]. This was first found in the study of massive integrable models[69, 70, 71] and the corresponding phenomena have been revisited several times[72, 73, 74, 66] with a connection to the Moore-Seiberg data[75, 76, 77] or the Li-Haldane conjecture[78]. It should be pointed out that the topological entanglement entropy[79, 80] and entanglement spectrum have a close connection to such observations, by applying the analysis of entanglement surfaces[81] to the TQFTs.
Even in these elementary algebraic structures, subrings and ring homomorphisms, we demonstrate that the appearance of noninteger coefficients is inevitable by developing the approach in [48]. However, we strongly believe that it will be possible to formulate the arguments in this work by using some generalized category theories, such as premodular fusion categories with suitable generalizations. This research direction has been still under development, but our arguments will provide clues for constructing such extended categories by requiring their compatibility with abstract algebra or linear algebra.
For readers in high-energy physics, we demonstrate that these extended category theories will play a fundamental role in studying the quantum phase transitions related to the Higgs mechanism[82] in the recent literature[83] (see Sec. IV). We clarify straightforward, but nontrivial relationships between the tensor structure of CFTs and the homomorphism in a general way. We note that the corresponding phenomena have been studied by several distinct research communities in different names, color confinement[84], $\lambda$-deformations[85, 86], and coupled-wire construction[87, 88, 89] (or related junction or ladder models) in condensed matter theory[90, 91, 92, 93, 94].
In this work, we study the massless and massive RG flows of two-dimensional CFTs, focusing on the conserved charges forming the fusion rings satisfying the NIM-reps. More phenomenologically, we study the classification of (non)unitary CFTs describing local (pseudo-)Hermitian systems and their quantum phase transition. Most part of the techincal details in this work only require the knowledge of linear algebra, and our method is applicable to higher-dimensional systems in principle. The rest of the manuscript is organized as follows. Sec. II is a short introduction to the pseudo-Hermitian systems. We introduce the fundamental properties of (pseudo-)Hermitian systems, which should be correct for the corresponding (non)unitary CFT. We note that this aspect has rarely been discussed in the literature on QFTs. In Sec. III, we revisit the fusion ring symmetry formed by the Verlinde line operators of the nonunitary CFTs from the modern view of pseudo-Hermitian systems. The generalization of quantum dimensions, which are less common in literature, is introduced, and its phenomenologies are studied in this section. We also study the general classification of gapped phase preserving Fibonacci fusion ring symmetry under a few assumptions based on the pioneering works on massive RG[61] and defect and boundary CFTs[95]. Sec. IV is the main part of this work, and we introduce a general method for constructing and analyzing the homomorphism between fusion rings. We demonstrate the fundamental importance of introducing the generalization of quantum dimensions. We also clarify the relation between the massless RG and coset construction or level-rank duality, which appeared commonly and has been studied (to some extent heuristically) in the literature. The discussion in this section also provides a straightforward but general understanding of the transitions in [83]. In the successive section, Sec. V, we study the massless RGs of several known examples. In Sec. VI, we introduce the anomaly classification for non-group-like objects appearing in the homomorphism by generalizing the notion of the integer spin simple current conditions (or equivalent anomaly-free conditions) in [96, 97, 98, 99, 66] (See also the earlier works [100, 101, 102, 103] and recent related works [104, 105]). Sec. VII is the concluding remarks of this work. In the Appendix, we comment on related research directions for the application of our methods to more general settings.
II Pseudo-Hermitian and linear dual basis
In this section, we introduce some basic aspects of pseudo-Hermitian systems[14, 15, 16] and several assumptions relevant to the successive sections. For a general review of pseudo-Hermitian systems, we note recent reviews[1, 2] in condensed matter and an older review [5] which has a close connection to the development of CFTs. We also note that related discussions have been summarized concisely in a recent work [106] with applications to the scattering theory.
First, we introduce the following data:
  • The Hamiltonian $H$ has real eigenvalues $\{E_{\lambda}\}$ labelled by $\{\lambda\}$.
  • Right energy eigenstates are $\{|\lambda\rangle\}$.
  • Linear dual $\{\widetilde{\langle\lambda|}\}$ of the right energy eigenstates satisfies $\widetilde{\langle\lambda|}|\eta\rangle=\delta_{\lambda,\eta}$ and forms left energy eigenstates.
where $\delta$ is the Kronecker delta. In other words, just by replacing the complex conjugation with the linear dual, the algebraic description of a general model is the same for Hermitian systems in our settings. To obtain the above properties of the given Hamiltonian, one needs to introduce more respective assumptions, such as $PT$ symmetry[17, 107], but the discussions will become more respective to introduce such assumptions. Hence, we assume the above properties [^4] We thank Naomichi Hatano for clarifying these points.. We also itemize the aspects that are unusual compared with Hermitian systems.
  • Hamiltonian $H$ does not satisfy the Hermiticity $H=H^{\dagger}$ in general.
  • Linear dual $\{\widetilde{\langle\alpha|}\}$ does not satisfy the relation $\widetilde{\langle\alpha|}=(|\alpha\rangle)^{\dagger}$ in general.
In other words, a Hermitian system is a special case of pseudo-Hermitian systems satisfying $H=H^{\dagger}$ and $\widetilde{\langle\alpha|}=(|\alpha\rangle)^{\dagger}$. In the successive sections, we demonstrate their interpretation in two-dimensional unitary and nonunitary CFTs  [^5] Because the nonunitary CFT we study in this section has real eigenenergy, it might be more precise to call them pseudo-unitary CFT or pseudo-Hermitian CFTs..
III Verlinde line operator and algebraic generalized quantum dimension
In this section, we revisit the structure of the Verlinde line operators with a formalism applicable to nonunitary conformal field theory corresponding to pseudo-Hermitian systems. We note [110, 111] as general references for CFTs, and [112, 113] as references for BCFTs.
For simplicity, we concentrate our attention on the $A$-type diagonal model. First, we assume the modular invariant as follows,
$Z=\sum_{\alpha}\chi_{\alpha}(t)\overline{\chi_{\alpha}}(\overline{t})$ (1)
where $\alpha$ is the label of primary fields and $t$ ($\overline{t}$) is the chiral (or antichiral) modular parameter and $\chi$($\overline{\chi}$) is the corresponding chiral (antichiral) character [^6] We have used the symbol $t$ which is less common than $\tau$ to represent the modular parameter for a notational reason. In the successive discussion, we use $\tau$ for the Fibonacci anyon, satisfying the fusion rule $\tau\times\tau=I+\tau$.. For latter use, we introduce the following modular $T$ and $S$ transformations,
$T:t$ $\rightarrow t+1,$ (2)
$S:t$ $\rightarrow\frac{-1}{t}.$ (3)
Phenomenologically, the modular $T$ properties characterize the locality of the systems and the modular $S$ properties characterize the high-low temperature duality of the systems via the open-closed duality[115].
By assuming the pseudo-Hermiticity of the Hamiltonian, one can express the chiral characters as follows,
$\chi_{\alpha}(t)=\sum_{M}\widetilde{\langle\alpha,M|}e^{2\pi it(L_{0}-c/12)}|\alpha,M\rangle$ (4)
where $L_{0}$ is the Virasoro generator forming the chiral part of the QFT Hamiltonian and $M$ is the label of descendant fields. For the latter use, we introduce the Virasoro algebra $\{L_{m}\}_{m\in\mathbb{Z}}$,
$L_{m},L_{n}]=(m-n)L_{m+n}+\frac{c}{12}(m^{3}-m)\delta_{m,-n}$ (5)
where $\mathbb{Z}$ is the set of integer numbers. Corresponding to the eigenvalue of $L_{0}$, and the primary field $\alpha$ has conformal dimension $h_{\alpha}$. By applying the $\{L\}_{m\geq-1}$ recursively to the primary states, one can obtain the corresponding descendant states denoted as $|\alpha,M\rangle$.
By identifying $\lambda=(\alpha,M)$, the states $\{|\alpha,M\rangle\}=\{|\lambda\rangle\}$ are in the scope of the general framework of pseudo-Hermitian systems. Hence, corresponding to this representation, we introduce the following projection operators,
$P_{\alpha,M}=|\alpha,M\rangle\widetilde{\langle\alpha,M|}$ (6)
One can implement corresponding representations for antichiral fields. Then, the following Verlinde line operator[39] can be written as follows,
$\mathcal{Q}_{\alpha}=\sum_{\beta,M,\overline{M}}\frac{S_{\alpha,\beta}}{S_{I,\beta}}P_{\beta,M}\overline{P}_{\beta,\overline{M}},$ (7)
where $S$ is the modular $S$ matrix defined by $\chi_{\alpha}(-1/t)=\sum_{\beta}S_{\alpha,\beta}\chi_{\beta}(t)$ and II is the identity operator corresponding to the vacuum. One can interprete the Verlinde operator as an exact symmetry of the CFT Hamiltonian because of the representation by projections, i.e., $\mathcal{Q}_{\alpha},H_{\text{CFT}}]=0$. This is a benefit of this quantum Hamiltonian method, and the set of symmetry operators $\{\mathcal{Q}_{\alpha}\}$ forms a ring over $\mathbb{C}$ by definition. One can easily detect the noninvertible conserved charges by identifying the zero of the modular-$S$ matrix, and this provides constraints of the phase diagram when studying the massive flows preserving the corresponding symmetries[66].
For latter use, we introduce the Verlinde formula [116],
$N^{\gamma}_{\alpha,\beta}=\sum_{\delta}\frac{S_{\alpha,\delta}S_{\beta,\delta}\overline{S_{\delta,\gamma}}}{S_{I,\delta}}$ (8)
where $N$ is the fusion matrix defined by the fusion rule $\alpha\times\beta=\sum_{\gamma}N^{\gamma}_{\alpha,\beta}\gamma$. Applying the Verlinde formula, one can obtain the following remarkable relation,
$\mathcal{Q}_{\alpha}\times\mathcal{Q}_{\beta}=\sum_{\gamma}N^{\gamma}_{\alpha,\beta}\mathcal{Q}_{\gamma}$ (9)
where $N$ is the nonnegative integer matrix and the symbol $\times$ represents the matrix multiplication. Hence, one can say that the Verlinde line is a typical object satisfying the algebraic relation between anyons in a CFT, called the fusion ring. This algebraic aspect has captured attentions in the field in the term, symmetry topological field theory (SymTFT)[117, 118, 119], by combining these algebraic data and Moore-Seiberg data[75, 76, 77]. We also note the related earlier work [120] which contains the basic idea of the sandwich construction in the recent terminolgy[121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 66, 131]. We stress that the fusion ring $\{\mathcal{Q}_{\alpha}\}$ is defined as a ring over $\mathbb{C}$. More phenomenologically, $\{\mathcal{Q}_{\alpha}\}$ should be treated as a set of conserved charges that has a close connection to topological defects forming a fusion category. Historically, this interpretation of conserved charges has appeared in [39], to demonstrate the physical realization of fusion category structure studied in earlier mathematical literature. In other words, in those days, conserved charges realizing rings over $\mathbb{C}$ were thought of as more physical than category theory. We also note that the distinction between conserved charges and topological defects plays a role in studying corresponding lattice realizations[132, 133, 134, 135, 136, 137].
We also note that the noninvertible symmetry operator defined by linear algebra (or zero of the modular $S$ matrix in our case) can be different from the “noninvertible symmetry" defined in category theories [^7] For example, the Fibonacci anyon $\tau$ with the fusion rule $\tau\times\tau=I+\tau$ generates invertible symmetry $\mathcal{Q}_{\tau}$ under the relation $\mathcal{Q}_{\tau}(\mathcal{Q}_{\tau}-\mathcal{Q}_{I})=\mathcal{Q}_{I}$.. In other words, there can exist categorically noninvertible symmetry operators that are invertible in linear algebra. However, when studying quantum Hamiltonian formalism based on linear algebra, the definition in linear algebra is more dominant when studying the RGs. We believe that there exists some categorical formalisms consistent with the observations in linear algebra, but they seem less common in the physics literature. To avoid confusion, we notify the “noninvertible symmetry" in category theory as categorical noninvertible symmetry or non-group-like symmetry. It also seems reasonable to call the corresponding objects nonabelian symmetry when considering its connection to nonabelian anyons in physics, but the term “nonabelian” also appears in group theory and the corresponding group rings in a different meaning.
In this paper, we propose the following quantity as a generalization of the quantum dimension, algebraic generalized quantum dimension,
$q_{\alpha,(a)}=\frac{S_{\alpha,a}}{S_{I,a}}$ (10)
where $\alpha$ represent type of symmetry or anyons, $a$ represents sector. (We note that our generalization is different from the generalized quantum dimension introduced in the study of representation theory of Grassmannian[139].) By choosing $a$ as the label of the lowest energy state $o$ or the effective vacuum, it is known that the above become positive numbers in unitary CFTs and a wide class of nonunitary CFTs[140, 141]. In a unitary CFT, the relation $o=I$ holds, but in a nonunitary CFT, this does not hold. This swapping of the role of the exact vacuum $I$ and the effective vacuum $o$ is called Galois shuffle[140] and played a significant role in proving modular noninvariance of gapless fractional quantum Hall states[142, 143, 144, 98, 66]. We also note that the quantity $q_{\alpha,(a)}$ has already appeared in a few mathematical studies [145, 146], or as the coefficients of the Verlinde operators[39], as we revisited Eq. (7). However, its implications for RG flows have not been studied to our knowledge. One can also introduce the antichiral analog of the algebraic generalized quantum dimensions $\overline{q_{\alpha,(a)}}$, and the distinction coming from the chirality plays a fundamental role in studying more general models, such as $Z_{N}$ symmetric models[147, 66, 48, 131]. However, for simplicity, we mainly focus on the $Z_{2}$ symmetric cases, and focus on the chiral structures in theories. Problems resulting from the chiralities of theories are discussed in Sec. VI.
The appearance of this number can be reproduced straightforwardly, by inserting Eq.(7) to Eq. (9) and concentrating attention on the coefficient before the projection labeled by $a$, $P_{a}$. Hence, the following relation holds.
$q_{\alpha,(a)}\times q_{\beta,(a)}=\sum_{\gamma}N^{\gamma}_{\alpha,\beta}q_{\gamma,(a)}$ (11)
In other words, this is an entropic formula representing a conservation of degrees of freedom through the fusions. More mathematically, the quantum dimension can be interpreted as a mapping between the fusion ring $\mathbf{A}$ to $\mathbb{C}$
$\begin{split}d_{(a)}:\ \mathbf{A}&\rightarrow\mathbb{C},\\ d_{(a)}(\alpha)&=q_{\alpha,(a)}.\end{split}$ (12)
As far as we know, this algebraic interpretation of quantum dimension appeared earlier literature[148, 149, 145, 150] (based on the representation theory), but its phenomenological (or mathematical) implications have not been studied. In the successive discussions, we study their implications for the algebraic structures, modules, and ideals, which provide essential information in classifying the corresponding RGs and TQFTs.
III.1 (smeared) boundary conformal field theory and entropy
In this subsection, we propose the form of boundary conformal field theories (BCFTs) in a way applicable also to the smeared BCFTs[61]. The Cardy’s states $|\alpha\rangle$ are given as follows[151],
$|\alpha\rangle=\sum_{\beta}\frac{S_{\alpha,\beta}}{\sqrt{S_{I,\beta}}}|\beta\rangle\rangle$ (13)
Where the right-hand side is the corresponding Ishibashi states[152]. Corresponding to this set of Cardy’s states, one can realize the corresponding linear dual of Cardy’s states as follows,
$\widetilde{\langle\alpha|}=\sum_{\beta}\frac{\overline{S_{\alpha,\beta}}}{\sqrt{S_{I,\beta}}}\widetilde{\langle\langle\beta|}$ (14)
As can be seen from this expression, Cardy’s state is defined by the linear dual, not the complex conjugate. It should be remarked that because we have only modified the definition of the bra and ket to those based on the linear dual, the existing arguments on BCFT is applicable to our case. From this basis, one can obtain the following amplitude of Ishibashi states,
$\widetilde{\langle\langle\alpha|}e^{\pi it(L_{0}+\overline{L_{0}}-c/12)}|\beta\rangle\rangle=\delta_{\alpha,\beta}\chi_{\alpha}(t)$ (15)
Because of these definitions, one can obtain the following Cardy’s condition,
$\widetilde{\langle\alpha|}e^{\pi it(L_{0}+\overline{L_{0}}-c/12)}|\beta\rangle=\sum_{\gamma}N^{\gamma}_{\alpha,\beta}\chi_{\gamma}(-1/t),$ (16)
where we have applied the Verlinde formula for the NIM-rep[115]. Hence, the BCFT constructed from our formalism based on the analysis of pseudo-Hermitian systems perfectly produces the existing results on the BCFT spectrum realizable in the corresponding two-dimensional statistical models, such as RSOS models[153]. The BCFT will also describe the boundary phenomena of the corresponding anyonic chains[154, 155], but the corresponding numerical or combinatorial studies are limited. For further studies, we note related works on the defects in the anyonic models[156, 133, 134, 137].
The boundary entropy is defined by evaluating the dominant contributions of the limit $it\rightarrow$-$\infty$. From the open string basis, this is a high temperature limit and difficult to evaluate directly, but from the closed string basis, it can be evaluated easily by observing the coefficient before the lowest energy states. Hence, one can obtain the following dominant asymptotics,
$\widetilde{\langle\alpha|}e^{\pi it(L_{0}+\overline{L_{0}}-c/12)}|\beta\rangle\sim\frac{S_{\alpha,o}}{\sqrt{S_{I,o}}}\frac{S_{\beta,o}}{\sqrt{S_{I,o}}}e^{-\pi tic_{\text{eff}}/6}.$ (17)
where $c_{\text{eff}}=c-24h_{o}$ is the effective central charge taking a positive number. By taking the logarithm, one can obtain the following entropic formula,
$\frac{\pi itc_{\text{eff}}}{6}+\text{log}g_{\alpha,o}+\text{log}g_{\beta,o}$ (18)
where $g_{\alpha,o}=S_{\alpha,o}/\sqrt{S_{I,o}}$ is the so called boundary $g$ entropy[157]. We note that these values are positive also in many nonunitary CFTs[140, 141]. Hence, with some appropriate assumptions, such as $PT$-symmetry, we conjecture:
[…]
Compared with the literature introducing negative norm states[158], our discussions on nonunitary BCFTs are simpler. Moreover, one can straightforwardly apply the correspondence between BCFTs and massive RGs in [61], because we have defined both BCFTs and the symmetry operators in a consistent way, satisfying the following relation,
$\mathcal{Q}_{\alpha}|\beta\rangle=\sum_{\gamma}N^{\gamma}_{\alpha,\beta}|\gamma\rangle\left(=|\alpha\times\beta\rangle\right)$ (20)
One can check its consistency only by applying the Verlinde formula[116, 115]. This relation appeared evidently in [95] to our knowledge.
For readers intereted in the implication of these extended boundary states constructed from the application of the defects, Eq. (20) implies the Cardy states with large $g$-value are a consequence of fusing defects to the boundary, and this implements the realization of the general boundary conditions with large $g$-values or boundary degrees of freedom (or qubits)[159, 160]. One can interpret these boundary edge modes as a generalization of the Majorana edge modes[161, 162, 163] or protected edge modes in the symmetry-protected topological phases[164, 165]. For readers interested in these edge modes, we note works by the first author[166, 131] and a related work [167].
III.2 Examples of our quantum dimension
In this subsection, we demonstrate some concrete examples of our quantum dimensions in minimal models $M(p,q)$ labelled by integers $(p,q)$. Firstly, we consider the $M(2,5)$ model, which is the simplest non-unitary minimal model. The $M(2,5)$ model have two chiral primaries $\alpha=\{I,\tau\}$ whose conformal weight are $h_{I}=0,h_{\tau}=-1/5$. Hence, $\tau$ is the lowest energy state $\tau=o$. If we consider the naive quantum dimension i.e., we take $a=I$ for $q_{\alpha,(a)}$, we obtain negative quantum dimension,
$q_{I,{(I)}}=1,\;q_{\tau,{(I)}}=\frac{1-\sqrt{5}}{2}<0.$ (21)
On the other hand, if we take $a=\tau$, we have a positive quantum dimensions,
$q_{I,(\tau)}=1,\;q_{\tau,{(\tau)}}=\frac{\sqrt{5}+1}{2}>0.$ (22)
One can check the consistency of the quantum dimensions by replacing the symbols in the relation $\tau\times\tau=I+\tau$.
Next we consider $M(2,7)$. The $M(2,7)$ model has three chiral primaries $\{I,\Phi,\Psi\}$ whose conformal weights are $h_{I}=0,h_{\Phi}=-2/7,h_{\Psi}=-3/7$. For $a=I$, we again have negative quantum dimensions,
[…]
and we see $q_{\Phi,(I)},q_{\Psi,(I)}$ are negative.
On the other hand, when we take $a=\Psi=o$, we have positive quantum dimension,
[…]
It is worth noting that the definition of the quantum dimensions for nonunitary CFTs sometimes differs in the literature, but their mutual implications have not been studied to our knowledge (See [140] and [104], for example). We provide them in Sec. IV and study their implications in concrete exmples in Sec. V.
III.3 Possible gapped phase with Fibonacci fusion ring symmetry
In this subsection, we demonstrate the benefit of the combination of the smeared BCFT[61] and the Graham-Watts method[95] for the classification of gapped phases and their quantum states. We concentrate our attention on the Fibonacci fusion symmetry, but the same arguments can be applied to general models. We also note that we provide a complete list of the possible (or realizible) gapped phases, but the realizations in some specific lattice or QFT models need further specific argugements depending on the specific algebraic and analytical data. Our discussion can be regarded as a generalization of the arguments in [44, 99, 168], by formulating the problem as invariant linear subspace of the modules spanned by smeared BCFTs.
First, let us assume the existence of RG flow from a CFT to a gapped phase while preserving the Fibonacci fusion ring symmetry, $\{I,\tau\}$, with the following fusion rule,
$\tau\times\tau=I+\tau$ (25)
In Hamiltonian, the RG flow to gapped phase can be expressed as $H_{\text{CFT}}+H_{\text{pert}}$ where $H_{\text{pert}}$ preserve the Fibonacci symmetry $\mathcal{Q}_{\tau},H_{\text{pert}}]=0$. For simplicity, we concentrate our attention on $\{I,\tau\}$, but one can consider the RG flow preserving $\{I,\tau\}\otimes\mathbf{C}^{\prime}$ where $\mathbf{C}^{\prime}$ is some arbitrary symmetry operators [^8] However, to study a larger fusion ring symmetry $\mathbf{C}^{\prime\prime}(\supset\{I,\tau\})$, one needs to do the same analysis in the respective situation..
Next, we assume that the massive RG flow does not break extended symmetries such as W-symmetry, and the resultant smeared Cardy’s states are labelled by the primary fields of the original diagonal CFT. This condition itself is nontrivial and can be broken when the symmetry is anomalous[170, 171, 172], but the analysis is nontrivial even restricting our attention to such a well-organized theory. In this setting, the resultant gapped phase should be restricted to the linear subspace of the Hilbert space spanned by the following (smeared) Cardy’s states,
$\mathcal{H}=\{x|I\rangle+y|\tau\rangle:x,y\in\mathbb{C}\}$ (26)
where $x,y$ are arbitrary complex numbers. By definition, this Hilbert space is invariant under the action of the symmetry $\{\mathcal{Q}_{\alpha}\}_{\alpha}=\{\mathcal{Q}_{I},\mathcal{Q}_{\tau}\}$ implemented as follows,
$\mathcal{Q}_{\alpha}|\beta\rangle=|\alpha\times\beta\rangle.$ (27)
where $\beta$ is a linear sum of $I$ and $\tau$.
This action of symmetries on the boundary states has been established in the pioneering work by Graham-Watts[95], and its fundamental significance to the classification of gapped states has been studied in the works by the first author[66, 48]. It should be stressed that the amplitude of smeared boundary states is defined as the inner product of quantum states. Hence, the amplitude becomes a general complex number, whereas their algebraic structure is determined by the NIM-rep of the symmetry operators.
As a gapped phase, there exist two choices of one one-dimensional gapped phase,
$\mathcal{H}_{+}=\left\{x\left(|I\rangle+q_{\tau,+}|\tau\rangle\right):x\in\mathbb{C}\right\},$ (28)
$\mathcal{H}_{-}=\left\{x\left(|I\rangle+q_{\tau,-}|\tau\rangle\right):x\in\mathbb{C}\right\}$ (29)
where $q_{\tau,\pm}=(1\pm\sqrt{5})/2$.
The above two Hilbert spaces are determined by the following condition,
$\mathcal{Q}_{\tau}|\alpha\rangle=Q_{\tau,\alpha}|\alpha\rangle$ (30)
where $\alpha$ is a linear sum of $I$ and $\tau$ and $Q_{\tau}$ is the corresponding eigenvalue. Hence, the problem is reduced to an eigenvalue problem of the linear transformation. One can obtain the following equations expressing the eigenvalue and eigenvector,
$\mathcal{Q}_{\tau}\left(|I\rangle+q_{\tau,+}|\tau\rangle\right)$ $=q_{\tau,+}\left(|I\rangle+q_{\tau,+}|\tau\rangle\right)$ (31)
$\mathcal{Q}_{\tau}\left(|I\rangle+q_{\tau,-}|\tau\rangle\right)$ $=q_{\tau,-}\left(|I\rangle+q_{\tau,-}|\tau\rangle\right)$ (32)
It should be stressed that without changing the normalization, the symmetry operator $\mathcal{Q}_{\tau}$ produces an unusual factor $q_{\tau,\pm}$. By introducing the new normalization as $\mathcal{Q}_{\tau}/q_{\tau,\pm}$, one can obtain unbroken spontaneous symmetry $\{\mathcal{Q}_{I},\mathcal{Q}_{\tau}/q_{\tau,\pm}\}$. More generally, the notion of spontaneous symmetry breaking of a state $|\alpha\rangle$ for fusion ring symmetry should be defined up to a constant.
Interestingly, there exists a correspondence or duality between $\mathcal{H}_{+}$ and $\mathcal{H}_{-}$ by the following $Z_{2}$ operation[173],
$I$ $\rightarrow I,$ (33)
$\tau$ $\rightarrow I-\tau.$ (34)
In other words, the group automorphism for the fusion ring plays the role of duality $\mathcal{H}_{+}\leftrightarrow\mathcal{H}_{-}$. We note that the above operation cannot be implemented by the multiplication of the linear sum of $I$ and $\tau$ in the original theory. This is different from the usual duality induced from the duality objects, and it seems worth emphasizing. However, we point out that the group automorphisms have played a significant role in studying the construction of extended algebra and associated generalized Verlinde formula in the studies of BCFTs[170, 171, 172, 174, 175]. Application of the automorphisms for the classification of symmetry-protected topological phases can be seen in [176], and a more precise relation between [176] and our arguments is worth further study. More generally, the objects in an automorphism preserving the unbroken symmetry will exchange the possible gapped phases, and we will study such generalized duality (or $G$-ality) elsewhere.
Consequently, there exist three possible gapped phases,
One choice of two-dimensional SSB phase $\mathcal{H},$ (35)
Two choices of one-dimensional SUB phase $\mathcal{H}_{\pm}.$ (36)
where SSB represents spontaneous-symmetry-breaking and SUB represents symmetry-unbroken. By generalizing the notion of the spontaneous symmetry breaking to the fusion ring symmetry, the Fibonacci fusion ring symmetry is spontaneously broken only in $\mathcal{H}$. To our knowledge, this symmetry analysis for the resultant module of the gapped phase is absent in the literature, even when restricting our attention to the simplest model, the Fibonacci fusion ring. We also note that the algebraic objects $I+q_{\tau,\pm}\tau$ form one-dimensional ideals of the original theory, and these ideals induce ring homomorphisms to the one-dimensional ring $\{I\}$ isomorphic to $\mathbb{C}$ for the latter discussion.
More generally, one can obtain the possible Hilbert space of a gapped phase with symmetry $\mathbf{A}_{\text{ub}}(\subset\mathbf{A})$ by classifying $\mathbf{A}_{\text{ub}}$ invariant modules of $\mathbf{A}$ in this setting. One can also include symmetry-breaking boundary states and their modules (by studying the extended algebra $\mathbf{A}^{\text{ex}}(\supset\mathbf{A})$), and the problem itself can be described in the language of linear algebra. Unfortunately, the category theories in physics are insufficient to solve this problem in linear algebra at this stage. Hence, further studies on more general category theories are fundamental, and we believe that some kind of premodular fusion categories with proper generalizations will solve such classifications in general. We note the works by Kikuchi as related works[99, 168, 177, 178, 179, 180] and our discussions provide an elementary method for studying the consequences of the massive RGs only from the established knowledge of linear algebra. It should be noted that the above arguments can provide a series of possible gapped phases, but the problem of detecting which gapped phases are realized in a given model will require more specific arguments. For this purpose, further studies on the eigenvalues $Q_{\alpha_{\text{ub}},\beta}$ of the states $|\beta\rangle$ under the symmetry action $\mathcal{Q}_{\alpha_{\text{ub}}}\in\mathbf{A}_{\text{ub}}$ defined by the relation $\mathcal{Q}_{\alpha_{\text{ub}}}|\beta\rangle=Q_{\alpha_{\text{ub}},\beta}|\beta\rangle$ will be fundamental. Finally, we note several recent works studying or constructing gapped phases with non-group-like (or fusion category) symmetry[181, 182, 183, 136, 184, 185, 186, 187, 188, 66, 189, 190].
IV Massless flows and matching algebraic generalized quantum dimensions
The massless RG flow[57] (or the RG domain wall[36, 37]) can be understood as a projection from a ultraviolet (UV) theory to an infrared (IR) theory realizing the ring homomorphism between fusion rings[58, 59, 60]. One of the most surprising facts in [48] is that the ring homomorphism is not unique even when restricting our attention to the simplest flow $M(4,5)\rightarrow M(3,4)$. Hence, it is necessary to study further conditions to detect or characterize each ring homomorphism. In the following, we briefly review the discussions in [48] expressing the massless RGs or gapped domain walls[38] as ring homomorphisms and introduce the benefit of studying the algebraic generalized quantum dimensions in classifying the homomorphisms. The introduction of the algebraic generalized quantum dimension provides systematic data of preserved sectors $\mathbb{S}_{\rho,\times}$ in the following discussions, and this point is a significant progress from the past work [48]. As a reference for the mathematical discussion in this section, we note an established textbook[191]. We also note several recent works in the communities[147, 192, 193, 194, 195], but the exact method to determine the coefficients has not been studied sufficiently, except for the works on conformal interfaces[196, 197, 198, 199, 200]  [^9] The literature on conformal interface provides the corresponding analytical data, but they often require extensive analytical calculations.
First, we introduce the ring homomorphism from a UV fusion ring $\mathbf{A}$ to an IR fusion ring $\mathbf{A}^{\prime}$. In the subsequent discussions, we distinguish the UV and IR theories by a prime symbol “ ’". We note that they are equivalent to the transformation law of anyonic objects appearing in the functor of category theories or the RG or gapped domain wall in theoretical physics. The ring homomorphism $\rho:\mathbf{A}\rightarrow\mathbf{A}^{\prime}$ is a mapping satisfying the following conditions,
$\rho(\alpha+\beta)$ $=\rho(\alpha)+\rho(\beta),$ (37)
$\rho(\alpha\times\beta)$ $=\rho(\alpha)\times\rho(\beta)$ (38)
Especially when treating fusion rings with unit, $I$, one can obtain the relation $\rho(I)=I^{\prime}$. For simplicity, let us assume that the mapping is surjective. From the Noether’s second ring isomorphism theorem, one can identify the IR fusion ring $\mathbf{A^{\prime}}$ as $\mathbf{A^{\prime}}=\mathbf{A}/\text{Ker}\rho$, where $\mathbf{I}=\text{Ker}\rho$ is a kernel of $\rho$ satisfying $\rho(\mathbf{I})=\{0\}$. The set $\mathbf{I}$ forms an ideal of the UV theory $\mathbf{A}$. Inversely, the classification ideal provides all possible ring homomorphisms. An ideal $\mathbf{I}$ is a set satisfying the relation $\mathbf{A}\times\mathbf{I}=\mathbf{I}$. The remarkable property of the relation $\mathbf{A}\times\mathbf{I}=\mathbf{I}$ is that one can replace the ideal $\mathbf{I}$ to $\{0\}$. To our knowledge, the significance of the ideal in fusion rings has been first discussed in [202]  [^10] However, it should be kept in mind that this pioneering work appeared much earlier, before the appearance of the fusion category symmetry.
Corresponding to the fusion rules and algebraic generalized quantum dimensions, we propose the following as a classification criterion,
[…]
where $a$ is a label of primary fields and $\mathbb{S}_{\rho}$ is the set specifying the preserved sectors, as we demonstrate in the following discussions. In particular, by choosing $o\in\mathbb{S}$, the flow will correspond to an existing massless integrable RG flow respecting the $c_{\text{eff}}$-theorem.
By generalizing the arguments in [48], one can systematically generate the ideal corresponding to $\text{Ker}\rho$ respecting the condition $d_{(a)}(\text{Ker}\rho)=0$. First, let us consider one element $s_{1}=\sum C_{\alpha,1}\mathcal{Q}_{\alpha}$ with $\sum C_{\alpha,1}q_{\alpha,(a)}=0$ where $\{C_{\alpha,1}\}_{\alpha}$ is the set of coefficients. The following will form the ideal preserving sector $a$,
$|s_{1}|=\oplus_{\alpha}s_{1}\mathcal{Q}_{\alpha}$ (40)
where we have taken the direct sum as a linear space generated by the basis $\{s_{1}\mathcal{Q}_{\alpha}\}$. Mathematically, this is called the ideal generated by $s_{1}$ (This is often represented as $(s_{1})$ in literature. We use a less common notation because we apply the homomorphism $d$ to the ideal, $d(|s_{1}|)$, as we show below). Because of the compatibility of the summation and fusion products, one will obtain the relation
$d_{(a)}(|s_{1}|)=\{0\}.$ (41)
This relation also implies that the objects in $|s_{1}|$ are noninvertible by assuming the matrix representation of $\{\mathcal{Q}\}$ and their linear algebra. In a similar way, one can construct multiple $(a)$ sectors preserving homomorphism, and we denote the preserved sector as $\mathbb{S}_{\rho}$. Hence the following defines the $\mathbb{S}_{\rho}$ preserving ideal,
$d_{(a)}(|s_{1}|)=\{0\},\text{ for all $a\in\mathbb{S}_{\rho}$}.$ (42)
More generally, one can consider the joint or linear sum of the ideals $|s_{i}|$.
Hence one can construct the ring homomorphism preserving the sector $\mathbb{S}_{\rho}$ as follows,
$\rho_{(a)}:\mathbf{A}\rightarrow\mathbf{A}/\mathbf{I}_{\mathbb{S}_{\rho}}$ (43)
where $\mathbf{I}_{\mathbb{S}_{\rho}}$ is an ideal satisfying $d_{a}(\mathbf{I}_{\mathbb{S}_{\rho}})=\{0\}$ for all $a\in\mathbb{S}_{\rho}$. In other words, we can obtain the following phenomenologies,
[…]
We also note that one can obtain the corresponding symmetry-preserving domain walls[204] by combining the above discussions with the CFT/TQFT and sandwich construction.
Interestingly, the algebraic construction implies the existence of the massless flow, which preserves the general sector $(a)$. Even when considering a unitary model, this results in the remarkable consequence that the ground state is not necessarily preserved under the ring homomorphism. In other words, there will exist quantum phase transitions which preserve excited sectors labelled by $(a)$. To some extent, this kind of RGs are conjectural, and we note a few works [205, 206] in related research directions. However, when interpreting the phenomena as domain walls in $2+1$ dimensional TQFTs, the existence of such unusual homomorphisms is more reasonable (or even trivial).
When assuming the fusion algebraic structure of the UV and IR theories, the above analysis implies the following compatibility of quantum dimensions and homomorphisms for an arbitrary UV object $\alpha$,
$d_{(a)}(\alpha)=d_{(a^{\prime})}(\rho(\alpha))\text{ for all $(a,a^{\prime})\in\mathbb{S}_{\rho,\times}$}$ (45)
where we have used ’ to indicate the objects in the IR theories, and $\mathbb{S}^{\prime}_{\rho,\times}$ is the pair of the preserved sector from UV to IR. We also introduce the form of the ring homomorphism as follows,
$\rho(\alpha)=\sum_{\alpha}A^{\alpha^{\prime}}_{\alpha}\alpha^{\prime}$ (46)
with
$\rho(I)=I^{\prime}.$ (47)
By assuming the compatibility of the quantum dimensions and ring homomorphism, one can obtain the following relations,
$q_{\alpha,(a)}=\sum_{\alpha^{\prime}}A^{\alpha^{\prime}}_{\alpha}q_{\alpha^{\prime},(a^{\prime})},$ (48)
for the preserved sector $(a,a^{\prime})\in\mathbb{S}_{\rho,\times}$. Hence, if we assume that the number of primary fields in the UV and IR theories is $n$ and $n^{\prime}$ respectively, one will obtain n1n-1 sets of first-order linear equations with $(n-1)\times n^{\prime}$ variables for each fixed $(a)$. Hence, to obtain a solution uniquely, it is necessary to introduce $\mathbb{S}_{\rho,\times}$ with multiple objects in general. When the homomorphism is not surjective, the solution seems to represent a domain wall between TQFTs rather than a massless RG, whereas we do not provide an explicit proof in this manuscript. We also note that if one takes $(a,a^{\prime})\notin\mathbb{S}_{\rho,\times}$, the relation $q_{\alpha,(a)}=\sum_{\alpha^{\prime}}A^{\alpha^{\prime}}_{\alpha}q_{\alpha^{\prime},(a^{\prime})}$ does not hold. Hence, corresponding to the choice of algebraic generalized quantum dimension or the preserved reference sectors $\mathbb{S}_{\rho,\times}$ in the RG flows,
$\begin{split}\text{the preservation of quantum dimensions can be broken}.\end{split}$ (49)
This is a kind of $g$-theorem respecting the quantum symmetry operators and their actions on each quantum state. Phenomenologically, it is fundamental to check that the given homomorphism will preserve the vacuum sectors $(I)$ and $(I^{\prime})$ or the effective vacuum $(o)$ and $(o^{\prime})$ with a connection to $c$-theorem[18] or $c_{\text{eff}}$ theorem[19] respectively.
IV.1 Duality between massive and massless flows
In this subsection, we provide a more phenomenological explanation of RG domain walls and gapped domain walls based on the quantum physics of anyons. The discussions here are a variant of those in [48], emphasizing the aspects in quantum physics. First, let us consider anyon $\alpha$ in a two-dimensional CFT corresponding to a critical system with fusion symmetry algebra $\mathbf{A}$. In this setting, one can construct a measurement sending an ideal $\mathbf{I}\subset\mathbf{A}$ to zero only from a few calculations in linear algebra (see the corresponding discussions in [48]). One can obtain the IR theory $\mathbf{A}^{\prime}=\mathbf{A}/\mathbf{I}$ where “/” represents the quotient ring, and one can identify the measurement as application of the ring homomorphism $\rho:\mathbf{A}\rightarrow\mathbf{A}^{\prime}$. Hence, this measurement induces a quantum phase transition[207, 208, 209] and the anyon $\alpha\in\mathbf{A}$ will split to $\rho(\alpha)=\sum_{\alpha^{\prime}}A^{\alpha^{\prime}}_{\alpha}\alpha^{\prime}$ with $\alpha^{\prime}\in\mathbf{A}^{\prime}$. In other words, the massless RG flow is a particular type of measurement-induced quantum phase transition that projects a UV theory to an IR theory by reducing the fusion ring symmetries.
In the $2+1$ dimensional TQFTs, this mapping of anyons can be implemented by the application of domain walls[38, 210, 58, 47]. Interestingly, if the homomorphism $\rho$ preserves a subring $\mathbf{A}_{\text{ub}}\subset\mathbf{A}$ (or satisfying $\mathbf{A}_{\text{ub}}\cap\text{Ker}(\rho)=\{0\}$), the ideal $\mathbf{I}$ can be identified as $\mathbf{A}_{\text{ub}}$ invariant module when assuming the Moore-Seiberg data[75, 76, 77, 120] in the original ring $\mathbf{A}$, by identifying the labels of $\mathcal{H}$ to the algebraic objects in $\mathbf{A}$. In short, one can express this correspondence as an equivalence $\mathcal{H}=\mathbf{I}$ in the UV theory $\mathbf{A}$. In other words, there exists a general correspondence of condensable block $\mathbf{I}$ in a massless RG with unbroken symmetry $\mathbf{A}_{\text{ub}}$ and some module $\mathcal{H}$ of $\mathbf{A}_{\text{ub}}$ symmetry preserving massive RG flows (Fig.1).
Figure 1: Correspondence between massless and massive RG flows. Existence of an ideal $\mathbf{I}$ preserving $\mathbf{A}_{\text{ub}}$ implies the existence of the corresponding massive RG, by identifying $\mathbf{I}$ as the resultant module $\mathcal{H}$ in the $\mathbf{A}_{\text{ub}}$-preserving massive RG.
In the perturbative QFT formalism, this phenomenology itself can be explained more reasonably. First, let us introduce some perturbation $mH_{\text{pert}}$ preserving $\mathbf{A}_{\text{ub}}$ symmetry where $m$ is a coupling constant. In general, it is known that the sign of the coupling constant determines the property of RG flow, either massless or massive, whereas the unbroken symmetry is fixed to $\mathbf{A}_{\text{ub}}$. Hence, one can expect some kind of dual relation between the corresponding massless and massive RG flows, and the relation $\mathcal{H}=\mathbf{I}$ in the original UV theory $\mathbf{A}$ represents the expected duality. However, we also note that there can exist a module $\mathcal{H}$ which cannot be represented as an ideal in the UV theory $\mathbf{A}$. Hence, the existence of the massive RG flow cannot imply the corresponding existence of the massless RG flow in general, whereas the inverse is true for a series of models satisfying the Moore-Seiberg data[75, 76, 77, 120]. The discussions in this subsection seem to be related to gauging operations for non-group-like symmetries[211, 212, 213, 214, 215, 216], and the further clarification of their precise relationship is an interesting future problem.
IV.2 Tensor decomposition and gapped domain wall problem: A benefit of introducing algebraic generalized quantum dimension
In this subsection, we study the implications of the algebraic generalized quantum dimensions to the gapped or RG domain wall problem when the UV CFT $\mathbf{A}$ can be expressed as a tensor product of the domain wall $\mathbf{A^{\prime\prime}}_{\text{DW}}$ and IR CFT $\mathbf{A^{\prime}}$. For example, one can starightforwardly applies the argument in this section to the situation where the UV CFT, $G_{K}$ Wess-Zumino-Witten model[217, 218, 219], can be decomposed to the $G_{K}/H_{K}$ coset WZW model[220, 221, 222] and the $H_{K}$ WZW model, by identifying the domain wall as $G_{K}/H_{K}$ and the IR theory as $H_{K}$. We do not assume the chirality of objects and theories, and one can apply the arguments in this section to a more general setting by replacing the chirality of them, such as $\mathbf{A}\leftrightarrow\overline{\mathbf{A}}$. We discuss some involved problems resulted from the chirality in Sec. VI. We note references on related problems [147, 66, 48, 131].
First, we assume the following decomposition,
$\begin{split}\mathbf{A}&=\mathbf{A^{\prime}}\otimes\mathbf{A^{\prime\prime}}_{\text{DW}}\\ \alpha&=\sum_{\alpha^{\prime\prime}_{\text{DW}},\alpha^{\prime}}B^{\alpha}_{\alpha^{\prime},\alpha^{\prime\prime}_{\text{DW}}}\alpha^{\prime\prime}_{\text{DW}}\otimes\alpha^{\prime}\end{split}$ (50)
where $B$ is a nonnegative integer matrix realizing the tensor decomposition of the UV theory, $\alpha^{\prime}$ is an object in the IR theory, and $\alpha^{\prime\prime}_{\text{DW}}$ is an object in the domain wall theory. For the latter discussion, we stress that this decomposition implements the ring isomorphism $\rho$ between UV theory $\mathbf{A}$ and the tensor product of IR theory $\mathbf{A^{\prime}}$ by tracing out the domain wall theory $\mathbf{A^{\prime\prime}}_{\text{DW}}$.
Next, we assume the algebraic generalized quantum dimension is well defined for the domain wall $\mathbf{A^{\prime\prime}}_{\text{DW}}$ as follows,
$\begin{split}d_{(a^{\prime\prime}_{\text{DW}})}:\ \mathbf{A^{\prime\prime}}_{\text{DW}}&\rightarrow\mathbb{C}\\ d_{(a^{\prime\prime}_{\text{DW}})}(\alpha^{\prime\prime}_{\text{DW}})&=q_{\alpha^{\prime\prime}_{\text{DW}},(a^{\prime\prime}_{\text{DW}})}\end{split}$ (51)
Figure 2: Tensor decomposition of UV CFT and the resultant IR theory. The red color specifies the objects and operations in the theory at the domain wall and the blue color specifies the objects in the IR theory. By tracing out the theory at the domain wall by applying the homomorphism $d_{a^{\prime\prime}_{\text{DW}}}$ realizing the algebraic generalized quantum dimensions, one can obtain the homomorphism $\rho:\mathbf{A}\rightarrow\mathbf{A^{\prime}}$.
By applying this homomorphism to the domain wall part of the object in $\mathbf{A^{\prime}}\otimes\mathbf{A^{\prime\prime}}_{\text{DW}}$, one can implement the following ring homomorphism straightforwardly,
$\rho_{a^{\prime\prime}_{\text{DW}}}$ $=d_{a^{\prime\prime}_{\text{DW}}}\otimes I^{\prime}:\mathbf{A}\rightarrow\mathbf{A^{\prime}},$ (52)
$\rho_{a^{\prime\prime}_{\text{DW}}}(\alpha)$ $=\sum_{\alpha^{\prime\prime}_{\text{DW}},\alpha^{\prime}}B^{\alpha}_{\alpha^{\prime},\alpha^{\prime\prime}_{\text{DW}}}q_{\alpha_{\text{DW}},(a^{\prime\prime}_{\text{DW}})}\otimes\alpha^{\prime}$ (53)
Hence, this representation determines the coefficient of the homomorphism $A$ as follows,
$A^{\alpha}_{\alpha^{\prime}}=\sum_{\alpha^{\prime\prime}_{\text{DW}}}B^{\alpha}_{\alpha^{\prime},\alpha^{\prime\prime}_{\text{DW}}}q_{\alpha_{\text{DW}},(a^{\prime\prime}_{\text{DW}})}$ (54)
In other words, the multiplicity of the solutions corresponds to the choice of the domain wall particle $a^{\prime\prime}_{\text{DW}}$. Next section, we study homomorphisms between theories where the domain wall theories $\mathbf{A}^{\prime\prime}_{\text{DW}}$ are not known, but the existence of the corresponding massless RGs has been established in the literature.
Before moving into the next section, we note that the correspondence between the massless RGs and coset CFTs has already been studied in the literature with a connection to the color confinement [84]. There exists more general studies on related integrable deformations, called $\lambda$-deformation or current-current deformation[85, 86]. We note a recent review for further literature [223]. More recently, related domain walls and coset structure have been studied in [224, 83] by applying anyon condensation and a related mathematical approach[225, 226]. For more historical aspects, we note the corresponding remarks in the works by the first author[66, 48].
V Examples of ring homomorphism in unitary and nonunitary models
Before moving into the detailed studies, we note the ring isomorphism of the Fibonacci fusion rule. For readers interested in existing RG flows of nonunitary CFTs, we note several earlier references[227, 228, 229] and recent ones[230, 104]. Related structures have also been observed in the hierarchical structure of fractional quantum Hall states[231, 232, 233, 234, 235].
The Fibonacci fusion rule is formed by the two objects $\{I,\tau\}$ satisfying the following relations,
$\tau\times\tau=I+\tau,$ (55)
where II is the identity operator again. The ring isomorphism is necessary to satisfy the following relations,
$\rho(I)$ $=I,$ (56)
$\rho(\tau)$ $=A_{\tau}^{I}I+A_{\tau}^{\tau}\tau$ (57)
where $A_{\tau}^{I},A_{\tau}^{\tau}$ is constants satisying the requirement of ring isomorphism,
$\rho(\tau\times\tau)=\rho(\tau)\times\rho(\tau)$ (58)
We also assume $A_{\tau}\neq 0$ to obtain the surjective homomorphism. Then one obtains the following relations,
$A_{0}^{2}+A_{\tau}^{2}$ $=A_{0}+1,$ (59)
$2A_{0}A_{\tau}+A_{\tau}^{2}$ $=A_{\tau}$ (60)
Hence, one can obtain the following two solutions,
$(A_{0},A_{\tau})=(0,1),(1,-1)$ (61)
The solution $(A_{0},A_{\tau})=(0,1)$ corresponds to the trivial ring isomorphism and $(A_{0},A_{\tau})=(1,-1)$ is the nontrivial one. As far as we know, the ring isomorphism corresponding to the latter first appeared in [173]. Remarkably, this ring isomorphism combined with the quantum dimension satisfies the following relation,
$d_{(0)}(\rho(\tau))=1-q_{\tau,(\tau)}$ (62)
or equivalently,
$d_{(\tau)}(\rho(\tau))=1-q_{\tau,(0)}.$ (63)
One can confirm these relations by studying the $M(2,5)$ minimal model, known as Yang-Lee CFT. More phenomenologically, this changes the quantum dimensions to its dual of the characterstic polynomial equation $x^{2}=x+1$ obtained from the fusion rule $\tau\times\tau=I+\tau$. Because the corresponding ideal is trivial, only such an exchange is permitted. We also note that there exist two non-surjective ring homomorphisms,
$(A_{0},A_{\tau})=\left(\frac{1\pm\sqrt{5}}{2},0\right)$ (64)
for the latter use.
V.1 $M(4,5)$ to $M(3,4)$
First, we revisit the ring homomorphism in unitary CFTs from $M(4,5)$ to $M(3,4)$ with emphasis on the quantum dimensions[48]. The $M(4,5)$ model or tricritical Ising CFT can be represented as a tensor product of Ising and Fibonacci fusion ring, $\{I,\tau\}\otimes\{I,\psi,\sigma\}$. Following the analysis in [48], one can represent the homomorphism as follows,
$\rho(\{I,\psi,\sigma\})$ $=\{I^{\prime},\psi^{\prime},\sigma^{\prime}\},$ (65)
$\rho(\tau)$ $=A_{I^{\prime}}I^{\prime}+A_{\psi^{\prime}}\psi^{\prime}$ (66)
where we have assumed that the homomorphism $\rho$ acts on the Ising fusion ring part trivially, and the $Z_{2}$ anomaly is preserved. One can obtain the four solutions satisfying the above conditions,
$(A_{I^{\prime}},A_{\psi^{\prime}})=\left(\frac{1\pm\sqrt{5}}{2},0\right),\left(\frac{1}{2},\pm\frac{\sqrt{5}}{2}\right).$ (67)
Corresponding to the sign $\pm$, the resultant quantum dimensions $d_{(0^{\prime})}\left(\rho\left(\tau\right)\right)$ takes the form $q_{\tau,(0)}=\frac{1+\sqrt{5}}{2}$ or $q_{\tau,(\tau)}=\frac{1-\sqrt{5}}{2}$. Hence, the preserved sector, $I$ or $\tau$, can be different depending on the choice of homomorphism. We also remind that the conditions $A_{\psi^{\prime}}=0$ or $A_{\psi^{\prime}}\neq 0$ determine whether the fermion $Z_{2}$ parity is conserved or not. Hence, we conclude that the massless RG from $M(4,5)$ to $M(3,4)$ or the embedding of the theory $M(3,4)$ to $M(4,5)$ is determined by the shuffle structure of $\{I,\tau\}$ and fermion parity conditions.
V.2 $M(3,4)$ to $M(2,5)$
It is widely known that there exists a massless integrable flow between the Ising CFT and the Yang-Lee CFT, denoted as $M(3,4)\rightarrow M(2,5)$. The corresponding ring homomorphism has not been studied to our knowledge, whereas the related conformal interfaces have been studied in [236]. Hence, we provide the ring homomorphism that corresponds to the RG flow.
First, we fix the theories with nonnegative integer fusion coefficients for simplicity. The Ising CFT has the three chiral primary fields $\{I,\psi,\sigma\}$ with the following relations,
$\psi\times\psi$ $=I,$ (68)
$\psi\times\sigma$ $=\sigma,$ (69)
$\sigma\times\sigma$ $=I+\psi.$ (70)
where $\times$ represents the fusion product and II is the identitiy operator. The operator $\psi$ is called the (chiral) Majorana fermion field and generates $Z_{2}$ symmetry, and $\sigma$ is the (chiral) order field or spin field. In the literature on TQFT or TO, $\sigma$ is called an Ising anyon.
The $M(2,5)$ model has two objects $\{I,\tau\}$ that satisfy the Fibonacci fusion rule. To distinguish the UV and IR theory, we denote the IR objects as $\{I^{\prime},\tau^{\prime}\}$.
As one can see, the Ising CFT has three objects and the Yang-Lee CFT has two objects. Moreover, the Ising model has $Z_{2}$ group ring structure $\{I,\psi\}$. It is known that the $Z_{2}$ symmetry should be violated under the massless RG. Hence, the following is a candidate for the corresponding homomorphism,
$\rho(I)$ $=I^{\prime},$ (71)
$\rho(\psi)$ $=A_{I^{\prime}}^{\psi}I^{\prime}+A_{\tau^{\prime}}^{\psi}\tau^{\prime},$ (72)
$\rho(\sigma)$ $=A^{I^{\prime}}_{\sigma}I^{\prime}+A^{\tau^{\prime}}_{\sigma}\tau^{\prime},$ (73)
where $A_{0},A_{\tau}$ is constants satisying the requirement of ring homomorphism. For this case, the following are the sufficient conditions
$\rho(\psi\times\psi)=\rho(\psi)\times\rho(\psi),$ (74)
$\rho(\psi\times\sigma)=\rho(\psi)\times\rho(\sigma),$ (75)
$\rho(\sigma\times\sigma)=\rho(\sigma)\times\rho(\sigma),$ (76)
The first relation results in the following,
$(A_{\psi}^{I^{\prime}})^{2}+(A_{\psi}^{\tau^{\prime}})^{2}$ $=1,$ (77)
$2A_{\psi}^{I^{\prime}}A_{\psi}^{\tau^{\prime}}+(A_{\psi}^{\tau^{\prime}})^{2}$ $=0,$ (78)
Hence, the solution is,
$(A_{\psi}^{I^{\prime}},A_{\psi}^{\tau^{\prime}})=\pm(\frac{1}{\sqrt{5}},-\frac{2}{\sqrt{5}}),(\pm 1,0)$ (79)
The second condition results in the following relation,
$A_{\psi}^{I^{\prime}}A_{\sigma}^{I^{\prime}}+A_{\psi}^{\tau^{\prime}}A_{\sigma}^{\tau^{\prime}}$ $=A_{\sigma}^{I^{\prime}},$ (80)
$A_{\psi}^{I^{\prime}}A_{\sigma}^{\tau^{\prime}}+A_{\psi}^{\tau^{\prime}}A_{\sigma}^{I^{\prime}}+A_{\psi}^{\tau^{\prime}}A_{\sigma}^{\tau^{\prime}}$ $=A_{\sigma}^{\tau^{\prime}},$ (81)
The third condition results in the following equations,
$(A_{\sigma}^{I^{\prime}})^{2}+(A_{\sigma}^{\tau^{\prime}})^{2}$ $=1+A_{\psi}^{I^{\prime}},$ (82)
$2A_{\sigma}^{I^{\prime}}A_{\sigma}^{\tau}+(A_{\sigma}^{\tau^{\prime}})^{2}$ $=A_{\psi}^{\tau^{\prime}},$ (83)
For the solution $(A^{I^{\prime}}_{\psi},A^{\tau^{\prime}}_{\psi})=(0,-1)$, the second equation results in $(A^{I^{\prime}}_{\sigma},A^{\tau^{\prime}}_{\sigma})=(0,0)$. Hence, the solution does not provide a surjective homomorphism. For $(A^{I^{\prime}}_{\psi},A^{\tau^{\prime}}_{\psi})=(0,1)$, there exist two solutions,
$(A_{\sigma}^{I^{\prime}},A_{\sigma}^{\tau^{\prime}})=\pm(\frac{\sqrt{2}}{\sqrt{5}},-\frac{2\sqrt{2}}{\sqrt{5}})$ (84)
By applying the same analysis to , one can obtain,
$(A_{\sigma}^{I^{\prime}},A_{\sigma}^{\tau^{\prime}})=\pm(\frac{\sqrt{5}+1}{\sqrt{10}},-\frac{\sqrt{2}}{\sqrt{5}})$ (85)
for the case with $(A^{I^{\prime}}_{\psi},A^{\tau^{\prime}}_{\psi})=(1/\sqrt{5},-2/\sqrt{5})$, and
$(A_{\sigma}^{I^{\prime}},A_{\sigma}^{\tau^{\prime}})=\pm(\frac{\sqrt{5}-1}{\sqrt{10}},\frac{\sqrt{2}}{\sqrt{5}})$ (86)
fo the case with $(A^{I^{\prime}}_{\psi},A^{\tau^{\prime}}_{\psi})=(-1/\sqrt{5},2/\sqrt{5})$.
Hence, there exist six solutions by assigning the condition of ring homomorphisms, and it is necessary to interpret the solutions. For this purpose, we observe the algebraic generalized quantum dimensions and their compatibility under the homomorphism $\rho$ for the solution with $(A^{I^{\prime}}_{\psi},A^{\tau^{\prime}}_{\psi})=(0,1)$ for simplicity.
First, the sector $\rho(I)=\rho(\psi)=I^{\prime}$ does not produce nontrivial relations of quantum dimensions. Hence, we concentrate our attention on the relation $\rho(\sigma)=AI^{\prime}+B\tau^{\prime}$.
In general, one can calculate the quantum dimensions of $\rho(\sigma)$ as follows,
$d_{(a^{\prime})}(\rho(\sigma))=A+Bq_{\tau^{\prime},(a^{\prime})}$ (87)
When the UV vaccum flows to the sector $a^{\prime}$, the above implies the relation $A+Bq_{\tau^{\prime},(a)}=\sqrt{2}$.
Hence, one can interprete the solution $(A,B)=(\frac{\sqrt{2}}{\sqrt{5}},-\frac{2\sqrt{2}}{\sqrt{5}})$ as the vacuum, $(a,a^{\prime})=(I,I^{\prime})\in\mathbb{S}_{\rho,\times}$, preserving flow and the solution $(A,B)=(-\frac{\sqrt{2}}{\sqrt{5}},\frac{2\sqrt{2}}{\sqrt{5}})$ corresponds to the effective vacuum $(a,a^{\prime})=(I,\tau^{\prime})\in\mathbb{S}_{\rho,\times}$ preserving domain wall. Reflecing the choices of the preserved sector, the structure $\mathbb{S}_{\rho,\times}$ becomes $\mathbb{S}_{\rho,\times}=\{(I,I^{\prime})\}$ or $\mathbb{S}_{\rho,\times}=\{(I,\tau^{\prime})\}$ accordingly. Because of the $c_{\text{eff}}$-theorem for the $PT$-symmetric systems[19], the latter solution corresponds to the existing massless integrable flows. One can apply the same analysis to the other four solutions and their relation to the results in [236] is an interesting problem.
V.3 $M(3,5)$ to $M(2,5)$
In this section, we study the flow from the $M(3,5)$ minimal model to the $M(2,5)$ minimal model. First, the fusion rule of the $M(3,5)$ minimal model can be represented as $\{I,J\}\otimes\{I,\tau\}=\{I,J,\tau,J\tau\}$, where $J$ is the anomalous $Z_{2}$ simple current and $\tau$ is the Fibonacci anyon again. Then, by specifying the action of the ring homomorphism to the tensor decomposition, one can obtain the following homomorphism from the part $\{I,J\}$,
$\rho(I)$ $=I^{\prime}$ (88)
$\rho(J)$ $=A^{I^{\prime}}_{J}I^{\prime}+A^{\tau^{\prime}}_{J}\tau^{\prime}$ (89)
This results in the following relations,
$(A^{I^{\prime}}_{J})^{2}+(A^{\tau^{\prime}}_{J})^{2}$ $=1,$ (90)
$2A^{I^{\prime}}_{J}A^{\tau^{\prime}}_{J}+(A^{\tau^{\prime}}_{J})^{2}$ $=0,$ (91)
This results in the following surjective solutions, with $\rho:\{I,J\}\rightarrow\{I^{\prime},\tau^{\prime}\}$,
$(A^{I^{\prime}}_{J},A^{\tau^{\prime}}_{J})=\pm(\frac{1}{\sqrt{5}},-\frac{2}{\sqrt{5}})$ (93)
and nonsurjective ones, $\rho:\{I,J\}\rightarrow\{I^{\prime}\}$,
$(A^{I^{\prime}}_{J},A^{\tau^{\prime}}_{J})=(\pm 1,0),$ (94)
where the former two relations are nontrivial ring homomorphisms, and the rest correspond to the group homomorphisms $Z_{2}\rightarrow\{I\}$.
For the solution $(A^{I^{\prime}}_{J},A^{\tau^{\prime}}_{J})=\pm(\frac{1}{\sqrt{5}},-\frac{2}{\sqrt{5}})$, one can choose the action of homomorphism to $\{I,\tau\}$ as ring isomorphisms $\{I,\tau\}\rightarrow\{I,\tau^{\prime}\}$ or ring homomorphisms $\{I,\tau\}\rightarrow\{I^{\prime}\}$. Hence, there exists $2\times 4=8$ ring homomorphisms.
For the solution $(A^{I^{\prime}}_{J},A^{\tau^{\prime}}_{J})=\pm(1,0)$, it is necessary to introduce the ring isomorphism $\{I,\tau\}\rightarrow\{I,\tau^{\prime}\}$. Hence, there exist $2\times 2=4$ solutions. Totally, one can obtain $8+4=12$ homomorphisms. One of the remarkable points of the above homomorphism is the breaking of the $Z_{2}$ structure and the correspondence of the Fibonacci fusion ring sector. Interestingly, there exist solutions which break the $\{I,\tau\}$ but produce $\{I^{\prime},\tau^{\prime}\}$ from the sector $\{I,J\}$ for the case with $(A^{I^{\prime}}_{J},A^{\tau^{\prime}}_{J})=\pm(\frac{1}{\sqrt{5}},-\frac{2}{\sqrt{5}})$. In other words, the symmetry $\{I,\tau\}$ and the emergent symmetry $\{I^{\prime},\tau^{\prime}\}$ match accidentally or by coincidence in these homomorphisms [^11] The term, accidental, is different from that used to mean the small (or irrelevant) breaking of symmetry.. Further research on such accidental matching or coincidence of the symmetry subring in UV and the emergent symmetry subring at IR is worth further study.
V.4 $M(2,7)$ to $M(2,5)$
In this subsection, we consider the ring homomorphism from $M(2,7)$ to $M(2,5)$ as a more complicated example. The $M(2,7)$ minimal model has three chiral primaries $\{I,\Phi,\Psi\}$ with the fusion rule,
$\Phi\times\Psi$ $=\Phi+\Psi,$ (95)
$\Phi\times\Phi$ $=I+\Psi,$ (96)
$\Psi\times\Psi$ $=I+\Phi+\Psi.$ (97)
Following the analysis so far, we consider a ring homomorphism defined by
$\begin{split}\rho(I)&=I^{\prime},\\ \rho(\Phi)&=A_{\Phi}^{I^{\prime}}I^{\prime}+A_{\Phi}^{\tau^{\prime}}\tau^{\prime},\\ \rho(\Psi)&=A_{\Psi}^{I^{\prime}}I^{\prime}+A_{\Psi}^{\tau^{\prime}}\tau^{\prime}.\end{split}$ (98)
To preserve the fusion ring, we obtain six equations:
$A_{\Phi}^{I^{\prime}}A_{\Psi}^{I^{\prime}}+A_{\Phi}^{\tau^{\prime}}A_{\Psi}^{\tau^{\prime}}$ $=A_{\Phi}^{I^{\prime}}+A_{\Psi}^{I^{\prime}},$ (99)
$A_{\Phi}^{I^{\prime}}A_{\Psi}^{\tau^{\prime}}+A_{\Phi}^{\tau^{\prime}}A_{\Psi}^{I^{\prime}}+A_{\Phi}^{\tau^{\prime}}A_{\Psi}^{\tau^{\prime}}$ $=A_{\Phi}^{\tau^{\prime}}+A_{\Psi}^{\tau^{\prime}},$ (100)
$(A_{\Phi}^{I^{\prime}})^{2}+(A_{\Phi}^{\tau^{\prime}})^{2}$ $=1+A_{\Psi}^{I^{\prime}},$ (101)
$2A_{\Phi}^{I^{\prime}}A_{\Phi}^{\tau^{\prime}}+(A_{\Phi}^{\tau^{\prime}})^{2}$ $=A_{\Psi}^{\tau^{\prime}},$ (102)
$(A_{\Psi}^{I^{\prime}})^{2}+(A_{\Psi}^{\tau^{\prime}})^{2}$ $=1+A_{\Phi}^{I^{\prime}}+A_{\Psi}^{I^{\prime}},$ (103)
$2A_{\Psi}^{I^{\prime}}A_{\Psi}^{\tau^{\prime}}+(A_{\Psi}^{\tau^{\prime}})^{2}$ $=A_{\Phi}^{\tau^{\prime}}+A_{\Psi}^{\tau^{\prime}}.$ (104)
Since $\rho$ is surjective, we require $A_{\Phi}^{\tau^{\prime}}\neq 0$ and $A_{\Psi}^{\tau^{\prime}}\neq 0$. Using these conditions, the six equations can be reduced to a single polynomial equation for $A_{\Phi}^{I^{\prime}}$:
$\begin{split}&\quad 125(A_{\Phi}^{I^{\prime}})^{6}-250(A_{\Phi}^{I^{\prime}})^{5}-25(A_{\Phi}^{I^{\prime}})^{4}\\ &+225(A_{\Phi}^{I^{\prime}})^{3}-30(A_{\Phi}^{I^{\prime}})^{2}-45A_{\Phi}^{I^{\prime}}+1=0.\end{split}$ (105)
We confirm that this equation has six real solutions. For each solution, $A_{\Phi}^{\tau^{\prime}}$ is determined by
$A_{\Phi}^{\tau^{\prime}}=\frac{5(A_{\Phi}^{I^{\prime}})^{3}-5(A_{\Phi}^{I^{\prime}})^{2}+4A_{\Phi}^{I^{\prime}}-2}{-15(A_{\Phi}^{I^{\prime}})^{3}+10A_{\Phi}^{I^{\prime}}+3}.$ (106)
Similarly, $A_{\Psi}^{I^{\prime}}$ and $A_{\Psi}^{\tau^{\prime}}$ are determined by Eq. (101) and Eq. (102), respectively. Numerically, we obtain the following solutions:
$\begin{split}&(A_{\Phi}^{I^{\prime}},A_{\Phi}^{\tau^{\prime}},A_{\Psi}^{I^{\prime}},A_{\Psi}^{\tau^{\prime}})\\ =&(0.820079,0.606822,0.0407623,1.36352),\\ &(1.4269,-0.606822,1.40428,-1.36352),\\ &(-0.0226214,-0.756695,-0.426901,0.606822),\\ &(-0.40428,1.36352,1.02262,0.756695),\\ &(-0.779316,0.756695,0.179921,-0.606822),\\ &(0.959238,-1.36352,1.77932,-0.756695).\end{split}$ (107)
In conclusion, we find six solutions satisfying the homomorphism conditions. Interestingly, there exists only one positive solution $(A_{\Phi}^{I^{\prime}},A_{\Phi}^{\tau^{\prime}},A_{\Psi}^{I^{\prime}},A_{\Psi}^{\tau^{\prime}})=(0.820079,0.606822,0.0407623,1.36352)$. We conjecture that this positive solution corresponds to the existing massless RG flow, and further studies based on other methods are necessary for the confirmation of our prediction. From the analysis so far, we observe that these conditions lead to algebraic equations such as Eq. (105). It should be noted that the identification of the massless RGs as projections has been proposed in the study of $M(2,2k+1)$ models [238].
VI Spin-classification for non-group-like objects: (Half) integer spin conditions
Interestingly, there exists a nontrivial structure when restricting our attention to the chiral and antichiral property of the flow $M(4,5)\rightarrow M(3,4)$ preserving the Ising fusion ring, $\{I,\psi,\sigma\}\sim\{I^{\prime},\psi^{\prime},\sigma^{\prime}\}$. Assuming the $Z_{2}$ symmetry preservation, one can consider RG flows $\{I,\psi\}\rightarrow\{I,\psi^{\prime}\}$ and $\{I,\psi\}\rightarrow\{\overline{I}^{\prime},\overline{\psi}^{\prime}\}$ because these the $Z_{2}$ simple currents satisfying the $Z_{2}$ anomaly condition $h_{\psi}-h_{\psi^{\prime}}=0\ (\text{mod}.1)$ or $h_{\psi}+\overline{h}_{\overline{\psi}^{\prime}}=0\ (\text{mod}.1)$. This anomaly matching condition, a kind of ’t Hooft anomaly matching codition[239], is nothing but Lieb-Schultz-Mattis anomaly matching condition[240, 241, 242] with a close connection to the Haldane conjecture[243, 244, 245, 246, 247], and this condition results in the appearance of the general electron operators constructing the singlevalued wavefunction of the TQFT $\mathbf{A}\otimes\mathbf{A}^{\prime}$ or $\mathbf{A}\otimes\overline{\mathbf{A}}^{\prime}$ [204](or its $Z_{2}$ extension more precisely[248, 249, 250, 251, 252]). Hence, one will expect the same condition for the preserved structure $\{I,\psi,\sigma\}\sim\{I^{\prime},\psi^{\prime},\sigma^{\prime}\}$, but $\sigma$ and $\sigma^{\prime}$ only satisfies the relation $h_{\sigma}+h_{\sigma^{\prime}}=0$. Hence, the flow $\{I,\psi,\sigma\}\rightarrow\{I^{\prime},\psi^{\prime},\sigma^{\prime}\}$ only preserves $Z_{2}$ structure when applying the anomaly classification to the non-group like symmetry whereas the flow $\{I,\psi,\sigma\}\rightarrow\{\overline{I}^{\prime},\overline{\psi}^{\prime},\overline{\sigma}^{\prime}\}$ is an anomaly preserving flow. More categorically, these two flows correspond to the choice of the common connected etale algebra $\{I,\psi\}=\{I^{\prime},\psi^{\prime}\}$ or $\{I,\psi,\sigma\}=\{\overline{I}^{\prime},\overline{\psi}^{\prime},\overline{\sigma}^{\prime}\}$. Consequently, if one believes that the massless flow $M(4,5)\rightarrow M(3,4)$ preserves the Ising fusion ring structure in the anomaly-free way, the application of the corresponding domain wall should change the chirality of objects.
In other words, when introducing the notion of braiding and chiralities, the classification of the RGs and the corresponding domain walls in TQFTs becomes more nontrivial. However, one can still expect that the (half-)integer spin conditions, which appeared in the studies of the discrete torsion models [253, 254, 255, 100, 101, 102, 103], should govern the properties of the systems. The term is relatively less common, but this condition should be called (half-)integer spin non-simple current condition[225]. This condition has played a role in the studies of exceptional modular invariants, related heterotic string theories, and maverick coset CFTs[256, 257, 258, 259]. We note recent works[104, 105] investigating the corresponding structure and a related work[260].
For the flow $M(3,5)\rightarrow M(2,5)$, for example, the same phenomena will appear. The nontrivial point of the representation of the $M(3,5)=\{I,\psi\}\otimes\{I,\tau\}$ is that one can choose $(h_{\tau},h_{J\tau})=(1/5,-1/20)$ or $(h_{\tau},h_{J\tau})=(-1/20,1/5)$. Reflecting the conformal dimensions of the IR theories, $(h_{I^{\prime}},h_{\tau^{\prime}})=(0,-1/5)$, only the flow $\overline{M(3,5)}\rightarrow M(2,5)$ with $(h_{\tau},h_{J\tau})=(1/5,-1/20)$ satisfies the integer spin nonsimple current condtion $h_{\tau}+h_{\tau^{\prime}}=0$. The condition $h_{\tau}+h_{\tau^{\prime}}=0$ is the generalized version of the anomaly cancellation condition of nongroup-like structures in the UV and IR theories. More generally, one can express the anomaly-free conditions for the unbroken algebraic structures $a_{\text{ub}}\in\mathbf{A}_{\text{ub}}$, where an unbroken structure satisfies the relation $\rho(\mathbf{A}_{\text{ub}})=\mathbf{A}_{\text{ub}}$, as follows,
  • The anomaly matching flow, $\mathbf{A}\rightarrow\mathbf{A^{\prime}}$, satisfies $h_{a_{\text{ub}}}-h^{\prime}_{a^{\prime}_{\text{ub}}}\in\mathbb{Z}/2$ for all $a_{\text{ub}}\in\mathbf{A}_{\text{ub}}$.
  • The anomaly cancellation flow, $\overline{\mathbf{A}}\rightarrow\mathbf{A^{\prime}}$ satisfies $h_{a_{\text{ub}}}+h^{\prime}_{a^{\prime}_{\text{ub}}}\in\mathbb{Z}/2$ for all $a_{\text{ub}}\in\mathbf{A}_{\text{ub}}$.
where we have distinguished the IR objects by adding “’", as in the previous sections and $\mathbb{Z}/2$ represents the set of integers or half-integers. One can expect appearance of “electron operators" with the fusion rule $\mathbf{A}_{\text{ub}}$ which construct single-valued wavefunctions of the corresponding coupled TQFTs, $\mathbf{A}\otimes\mathbf{A^{\prime}}$ or $\overline{\mathbf{A}}\otimes\mathbf{A^{\prime}}$, but this research direction has been studied only in several cases, as we have remarked [256, 257, 258, 259]  [^12] Corresponding to this anomaly-free condition of the non-group like objects, there will exist intrinsically nonabelian anyonic TOs, or $\mathbf{A}_{\text{ub}}$-symmetry enriched TOs where the fundamental “electron operators” constructing the wavefunctions are nonabelian anyon belonging to $\mathbf{A}_{\text{ub}}\boxtimes\mathbf{A^{\prime}}_{\text{ub}}$ or $\overline{\mathbf{A}_{\text{ub}}}\boxtimes\mathbf{A^{\prime}}_{\text{ub}}$..
Before moving into the concluding remark, we note the unusual matching of the symmetry ring appearing in the flow $M(3,5)\rightarrow M(2,5)$. As we have demonstrated, there exist homomorphisms which break UV symmetry $\{I,\tau\}$ but (emergently) produce the IR symmetry $\{I^{\prime},\tau^{\prime}\}$ accidentally or by coincidence from the UV $Z_{2}$ group ring $\{I,J\}$. For the readers interested in such unusual emergent symmetry, we formulate the definition of the coincident emergent symmetry or coincidence of the UV-IR symmetries. First, let us assume subring of the UV fusion ring, $\mathbf{A}_{\text{coi}}\subset\mathbf{A}$ and there exists the IR subring $\mathbf{A^{\prime}}_{\text{coi}}\subset\mathbf{A^{\prime}}$ which are ring isomorphic to $\mathbf{A^{\prime}}$, i.e. $\mathbf{A}_{\text{coi}}=\mathbf{A^{\prime}}_{\text{coi}}$. One can identifies $\mathbf{A^{\prime}}_{\text{coi}}$ as coincident emergent symmetry when $\rho(\mathbf{A}_{\text{coi}})\neq\mathbf{A^{\prime}}_{\text{coi}}$ under a homomorphism $\rho$. In other words, the same structure in the UV and IR theories is produced under the action of $\rho$ to other structures $\mathbf{A}_{\text{ot}}\neq\mathbf{A}_{\text{coi}}$. There exist related discussions in [104], but the exact implications of this matching have not been studied to our knowledge. By combining the analysis on the anomaly-free condition in this section, one can find more nontrivial examples.
VII Conclusion
In this work, we have studied fusion ring symmetry of pseudo-Hermitian systems and their RG flows in a rigorous abstract algebraic formalism. The algebraic interpretation of the generalization of quantum dimensions plays the most fundamental role in our classifications, and these structures result in the appearance of (unusual) non-integer coefficients in the classifications. The notion of the spontaneous symmetry breaking of generalized symmetry for the gapped phase has been introduced in a linear algebraic way applicable to more general systems. Moreover, we provide a series of algebraic data of RG domain walls in CFTs or gapped domains in the corresponding TQFTs. We have explicitly demonstrated that the appearance of non-integer coefficients is inevitable when assuming the ring homomorphism structures, which are also fundamental for the categorical formulations of such domain walls. In other words, if one believes the RG domain walls or gapped domain walls are compatible with categorical formulations, it is necessary to introduce generalized category theories applicable to linear algebra (or intrinsically quantum systems). If one believes that the existing category theory without noninteger coefficients is sufficient, one cannot apply the notion of the functor and subcategory to the analysis of massless and massive RG flows. We note that this view has validity when studying defect and boundary RG flows, and this research direction is worth further study, whereas they are outside of the main scope of the present manuscript. Finally, we note that our method is applicable to higher-dimensional systems in principle and cite some recent references [262, 263, 192, 195].
VIII Acknowledgement
The authors thank Po-Yao Chang, Naomichi Hatano, and Hosho Katsura for sharing their knowledge on non-Hermitian systems. YF thanks Yuma Furuta and Shinichiro Yahagi for related collaborations. YF also thanks Sylvain Ribault and Ingo Runkel for the discussions on fusion rules and Jurgen Fuch and Yuji Tachikawa for the helpful discussions on the distinction between fusion categories and fusion rings. We thank Ingo Runkel and Gerard Watts for helpful comments and for checking some technical calculations in the manuscript. Also, we thank Tadashi Takayanagi for commenting on topological entanglement entropy and its importance. YF thanks the support from NCTS. TK is supported by Grant-in-Aid for JSPS Fellows No. 23KJ1315.
Appendix A Comments on conjectural structures in pseudo-Hermitian models
A.1 Comment on noncommutative fusion ring and magma
In this section, we comment on the noncommutative fusion ring and magma, outside of the scope of the main text. For a noncommutative fusion ring, there can exist matrix representations, but the underlying operations can be intertwining operations. This intertwining operation comes from degeneracies of the spectrum between different primary fields[39]. Hence, for example, a $Z_{N}$ symmetric model will naturally have this structure. Because of the noncommutativity, one cannot map the fusion ring to $\mathbb{C}$. Hence, quantum dimension will be defined in a different way as a ring homomorphism to some noncommutative number fields.
Next, we discuss the magma appearing in recent literature[264] [^13] More precisely, linear magma satisfying the pentagon relation.. In such models, the fusion rule can be nonassociative. Hence, there exists no matrix representation by identifying the fusion as matrix multiplication. However, by changing the definition of multiplication, it is possible to obtain a magma from the representation of matrices or linear operators. For example, commutator [A,B]=ABBA[A,B]=AB-BA and anticommutator $\{A,B\}=AB+BA$ are naive candidates which implement the fusion $\times$, by replacing “×\ $\times\$" to [,][\ ,\ ] or $\{\ .\ \}$. Even in this magma, the bilinearity of objects will survive, and this will simplify the calculations. We note a historical review for this research direction[266]. In this setting, the quantum dimension will be generalized to the mapping, magma homomorphism, to nonassociative number fields.
A.2 Phenomenology on entanglement cut: Naturality of complex spectrum
In the main texts, we studied two types of boundary phenomena: smeared BCFT appearing in a gapped phase or BCFT appearing in physical boundaries. In this subsection, we discuss some phenomenology on the entanglement surface[81, 267, 268].
First thing to note is that the entanglement spectrum in a pseudo-Hermitian system can be complex[269, 270, 271, 272]. However, following the discussion in the previous sections, this phenomenon seems unusual because the complex energy value cannot appear in the previous discussion. Moreover, there exist established studies on the universal behaviors of real entanglement entropy in the nonunitary CFTs by analyzing the corresponding two dimensional statistical models and their quantum group structure[273, 274], combined with the twist field method[275, 276, 277]. We also note [278] as a reference summarizing a variant of related entropies, such as those in [279, 280].
The most fundamental difference between our discussions and the phenomenology of the entanglement cut in recent studies is that the basis of the entanglement cut is not a linear dual basis. When keeping in mind this difference, the appearance of a complex spectrum is reasonable. Moreover, when assuming such an unusual basis, some condition of charge conservation will be broken, because we have chosen an unusual basis from the entanglement cut. Hence, some charged character will appear, and one of the most natural candidates is the nonspecialized character appearing in the corner-transfer-matrix (CTM) method[281, 282]. It is known that there exists a correspondence between CTM and the entanglement Hamiltonian, and the complex-valued CTM for a series of two-dimensional statistical models have been studied in [283, 284]. Hence, as in the Hermitian systems, the following correspondence should be true for the non-Hermitian systems,
[…]
For readers interested in the phenomenological understanding of the nonspecialized character, we note related works on the charged Cardy formula[285, 286, 287, 288]. It is worth mentioning that the nontriviality of the nonspecialized characters has already been remarked in the earlier works in defect and boundary CFTs[289, 39].
We also note that the appearance of the complex entanglement spectrum can be explained by using pseudo-entropy theories [290, 291, 292, 8] which implement the natural entropy under the final state projection. Moreover, the complex entanglement spectrum appears from the "double cone" regularization, which is one way of regulating UV divergence of QFT by introducing complex deformations [293, 294, 295]. In this setting, the entanglement Hamiltonian is non-Hermitian.
We remind that even in Hermitian systems, it has been reported that the Cardy state with the highest boundary entropy appears at the entanglement surfaces[296]. This seems unusual from the $g$-theorem[157], but one can understand this as a consequence of high-low temperature duality phenomenologically. The open-closed duality is a relation between high-temperature phenomena and low-temperature phenomena. In high-energy physics, this has been used to evaluate or estimate high-temperature or UV phenomena in those of IR. However, when considering the low-temperature physics realized in the ground states of a quantum lattice model, the opposite is used. Hence, more high-temperature (or UV-like) theories will appear at the entanglement surface when studying more low-temperature (or IR-like) theories. By combining this understanding with symmetry analysis of boundaries, one can expect the appearance of boundary states with large $g$-values (including symmetry-breaking branes[170, 171, 172] and the Graham-Watts(GW) states[95]) at the entanglement surfaces. The same reasoning will be true in non-Hermitian systems, but one needs to study the complex conjugate operation or linear dual operation more carefully.

### v2

@@
 One of the most fundamental ideas in the studies on non-Hermitian systems is the similarity transformation, which maps a pseudo-Hermitian system to the Hermitian system [14, 15, 16].
-Especially, when the system has the $PT$ symmetry, this has played an important role[17].
+Especially, when the system has the parity-time (PT) symmetry, this has played an important role[17].
 In quantum field theory, a kind of law of thermodynamics, called Zamolodchikov’s $c$-theorem[18] has been generalized to $c_{\text{eff}}$-theorem[19] for $PT$ symmetric models where $c$ is the central charge of a CFT and $c_{\text{eff}}$ is the effective central charge.
@@
 Moreover, it is known that the similarity transformation can change the locality of the system, and this is analogous to the gauge transformations[20].
-Hence, at this stage, a local pseudo-Hermitian system is more treatable than the corresponding nonlocal Hermitian system in QFT  [^1] We will study the symmetry topological field theories of nonlocal Hermitian systems constructed from local pseudo-Hermitian systems by the similarity transformations elsewhere..
+Hence, at this stage, a local pseudo-Hermitian system is more treatable than the corresponding nonlocal Hermitian system in QFT (For readers interested in the implications of the similarity transformation, we note the complementary work by the authors [21] and the references therein).
 In this sense, characterizations of pseudo-Hermitian systems and their quantum phase have still been under development in the QFT language.
@@
 Corresponding to the eigenvalue of $L_{0}$, and the primary field $\alpha$ has conformal dimension $h_{\alpha}$.
-By applying the $\{L\}_{m\geq-1}$ recursively to the primary states, one can obtain the corresponding descendant states denoted as $|\alpha,M\rangle$.
+By applying the $\{L\}_{m\leq-1}$ recursively to the primary states, one can obtain the corresponding descendant states denoted as $|\alpha,M\rangle$.
 By identifying $\lambda=(\alpha,M)$, the states $\{|\alpha,M\rangle\}=\{|\lambda\rangle\}$ are in the scope of the general framework of pseudo-Hermitian systems.
@@
 We also note the related earlier work [120] which contains the basic idea of the sandwich construction in the recent terminolgy[121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 66, 131].
+We also note that the application of SymTFTs to a class of nonequilibrium phenomena, Floquet systems, can be seen in [132, 133].
 We stress that the fusion ring $\{\mathcal{Q}_{\alpha}\}$ is defined as a ring over $\mathbb{C}$.
@@
 In other words, in those days, conserved charges realizing rings over $\mathbb{C}$ were thought of as more physical than category theory.
-We also note that the distinction between conserved charges and topological defects plays a role in studying corresponding lattice realizations[132, 133, 134, 135, 136, 137].
+Related to this aspect, it is known that the distinction between conserved charges and topological defects plays a role in studying corresponding lattice realizations[134, 135, 136, 137, 138, 139].
 We also note that the noninvertible symmetry operator defined by linear algebra (or zero of the modular $S$ matrix in our case) can be different from the “noninvertible symmetry" defined in category theories [^6] For example, the Fibonacci anyon $\tau$ with the fusion rule $\tau\times\tau=I+\tau$ generates invertible symmetry $\mathcal{Q}_{\tau}$ under the relation $\mathcal{Q}_{\tau}(\mathcal{Q}_{\tau}-\mathcal{Q}_{I})=\mathcal{Q}_{I}$..
@@
 However, when studying quantum Hamiltonian formalism based on linear algebra, the definition in linear algebra is more dominant when studying the RGs.
-We believe that there exists some categorical formalisms consistent with the observations in linear algebra, but they seem less common in the physics literature.
+We believe that there exists some categorical formalisms consistent with the observations in linear algebra, but they seem less common in the physics literature [^7] A $\mathbb{C}$-linear category will correspond to the linear algebra, but it is less common in the context of fusion rules.
 To avoid confusion, we notify the “noninvertible symmetry" in category theory as categorical noninvertible symmetry or non-group-like symmetry.
@@
 $q_{\alpha,(a)}=\frac{S_{\alpha,a}}{S_{I,a}}$ (10)
-where $\alpha$ represent type of symmetry or anyons, $a$ represents sector.
+where $\alpha$ represents type of symmetry or anyons, $a$ represents sector.
 (We note that our generalization is different from the generalized quantum dimension introduced in the study of representation theory of Grassmannian[142].) By choosing $a$ as the label of the lowest energy state $o$ or the effective vacuum, it is known that the above become positive numbers in unitary CFTs and a wide class of nonunitary CFTs[143, 144].
@@
 Problems resulting from the chiralities of theories are discussed in Sec.
-VI.
+VI with a possible connection to the Alice ring in literature[151, 152].
 The appearance of this number can be reproduced straightforwardly, by inserting Eq.(7) to Eq.
@@
 The boundary entropy is defined by evaluating the dominant contributions of the limit $it\rightarrow$-$\infty$.
-From the open string basis, this is a high temperature limit and difficult to evaluate directly, but from the closed string basis, it can be evaluated easily by observing the coefficient before the lowest energy states.
+From the open string basis, this is a high temperature limit and difficult to evaluate directly, but, from the closed string basis, it can be evaluated easily by observing the coefficient before the lowest energy states.
 Hence, one can obtain the following dominant asymptotics,
@@
 This relation also implies that the objects in $|s_{1}|$ are noninvertible by assuming the matrix representation of $\{\mathcal{Q}\}$ and their linear algebra.
-In a similar way, one can construct multiple $(a)$ sectors preserving homomorphism, and we denote the preserved sector as $\mathbb{S}_{\rho}$.
+In a similar way, one can construct homomorphisms preserving multiple $(a)$ sectors, and we denote the preserved sector as $\mathbb{S}_{\rho}$.
 Hence the following defines the $\mathbb{S}_{\rho}$ preserving ideal,
@@
 $d_{(a)}(\alpha)=d_{(a^{\prime})}(\rho(\alpha))\text{ for all $(a,a^{\prime})\in\mathbb{S}_{\rho,\times}$}$ (45)
-where we have used ’ to indicate the objects in the IR theories, and $\mathbb{S}^{\prime}_{\rho,\times}$ is the pair of the preserved sector from UV to IR.
+where we have used ’ to indicate the objects in the IR theories, and $\mathbb{S}_{\rho,\times}$ is the pair of the preserved sector from UV to IR.
 We also introduce the form of the ring homomorphism as follows,
@@
 $\rho(I)$ $=I^{\prime},$ (71)
-$\rho(\psi)$ $=A_{I^{\prime}}^{\psi}I^{\prime}+A_{\tau^{\prime}}^{\psi}\tau^{\prime},$ (72)
+$\rho(\psi)$ $=A^{I^{\prime}}_{\psi}I^{\prime}+A^{\tau^{\prime}}_{\psi}\tau^{\prime},$ (72)
 $\rho(\sigma)$ $=A^{I^{\prime}}_{\sigma}I^{\prime}+A^{\tau^{\prime}}_{\sigma}\tau^{\prime},$ (73)
@@
 Using these conditions, the six equations can be reduced to a single polynomial equation for $A_{\Phi}^{I^{\prime}}$:
-$\begin{split}&\quad 125(A_{\Phi}^{I^{\prime}})^{6}-250(A_{\Phi}^{I^{\prime}})^{5}-25(A_{\Phi}^{I^{\prime}})^{4}\\ &+225(A_{\Phi}^{I^{\prime}})^{3}-30(A_{\Phi}^{I^{\prime}})^{2}-45A_{\Phi}^{I^{\prime}}+1=0.\end{split}$ (105)
+$\begin{split}&\quad 125(A_{\Phi}^{I^{\prime}})^{6}-250(A_{\Phi}^{I^{\prime}})^{5}-25(A_{\Phi}^{I^{\prime}})^{4}\\ &+225(A_{\Phi}^{I^{\prime}})^{3}-30(A_{\Phi}^{I^{\prime}})^{2}-45A_{\Phi}^{I^{\prime}}-1=0.\end{split}$ (105)
 We confirm that this equation has six real solutions. For each solution, $A_{\Phi}^{\tau^{\prime}}$ is determined by
@@
 In conclusion, we find six solutions satisfying the homomorphism conditions.
-Interestingly, there exists only one positive solution $(A_{\Phi}^{I^{\prime}},A_{\Phi}^{\tau^{\prime}},A_{\Psi}^{I^{\prime}},A_{\Psi}^{\tau^{\prime}})=(0.820079,0.606822,0.0407623,1.36352)$.
+Interestingly, there exists only one positive solution $(A_{\Phi}^{I^{\prime}},A_{\Phi}^{\tau^{\prime}},A_{\Psi}^{I^{\prime}},A_{\Psi}^{\tau^{\prime}})=(0.820079,0.606822,0.0407623,1.36352)$, and the resultant homomorphism has the preserving sectors $\mathbb{S}_{\rho,\times}=\{(I,I^{\prime}),(o,o^{\prime})\}$.
 We conjecture that this positive solution corresponds to the existing massless RG flow, and further studies based on other methods are necessary for the confirmation of our prediction.
+Moreover, there exists a controversy between the perturbative expressions on the flow $M(2,7)\rightarrow M(2,5)$ in [243, 244] and those in [245], and this controversy will be explained by the nonuniqueness of the homomorphism in this subsection [^12] We thank Andrei Katsevich for notifying us the subtelties of $M(2,7)\rightarrow M(2,5)$ in the literature..
 From the analysis so far, we observe that these conditions lead to algebraic equations such as Eq.
@@
 Consequently, if one believes that the massless flow $M(4,5)\rightarrow M(3,4)$ preserves the Ising fusion ring structure in the anomaly-free way, the application of the corresponding domain wall should change the chirality of objects.
+Phenomenologically, the domain wall changing the chiralities seems to correspond to the Alice ring in literature[151, 152].
 In other words, when introducing the notion of braiding and chiralities, the classification of the RGs and the corresponding domain walls in TQFTs becomes more nontrivial.
@@
 YF also thanks Sylvain Ribault and Ingo Runkel for the discussions on fusion rules and Jurgen Fuch and Yuji Tachikawa for the helpful discussions on the distinction between fusion categories and fusion rings.
-We thank Ingo Runkel and Gerard Watts for helpful comments and for checking some technical calculations in the manuscript.
-Also, we thank Tadashi Takayanagi for commenting on topological entanglement entropy and its importance.
+We thank Hosho Katsura, Ingo Runkel and Gerard Watts for helpful comments and for checking some technical calculations in the manuscript.
+Also, we thank Tadashi Takayanagi for commenting on topological entanglement entropy and its importance, and Andrei Katsevich notifying us controvercies in the flow $M(2,7)\rightarrow M(2,5)$.
 YF thanks the support from NCTS.
@@
 Moreover, there exist established studies on the universal behaviors of real entanglement entropy in the nonunitary CFTs by analyzing the corresponding two dimensional statistical models and their quantum group structure[282, 283], combined with the twist field method[284, 285, 286].
-We also note [278] as a reference summarizing a variant of related entropies, such as those in [279, 280].
+We also note [287] as a reference summarizing a variant of related entropies, such as those in [288, 289] (for a recent analysis on symmetry resolved entanglement entropy[290, 291, 292, 293, 294, 295, 296, 297, 298] in nonunitary CFTs, see [299]).
 The most fundamental difference between our discussions and the phenomenology of the entanglement cut in recent studies is that the basis of the entanglement cut is not a linear dual basis.